Cream and Crimson Foods has a target capital structure of calling for 43.00 percent debt, 5.00 percent preferred stock, and 52.00 percent common equity (retained earnings plus common stock). Its before-tax cost of debt is 11.00 percent. The tax rate is 40.00%. Its cost of preferred stock is 11.70%. Its cost of common equity is 13.53%.

Find the WACC for Cream and Crimson Foods?
Answer #1

After-tax cost of debt=11*(1-tax rate)

=11*(1-0.4)=6.6%

WACC=Respective costs*Respective weight

=(0.43*6.6)+(0.05*11.7)+(0.52*13.53)

which is equal to

=10.4586%